You’ve observed the following returns on Yamauchi Corporation’s stock over the past five years: –29.7 percent, 16.8 percent, 36.6 percent, 3.9 percent, and 22.9 percent.

What was the variance of the returns over this period? (Do not round intermediate calculations and round your answer to 6 decimal places, e.g., 32.161616.)
Answer #1

Mean return=(-29.7%+16.8%+36.36%+3.9%+22.9%)/5=0.10052

Now, for each number, we need to subtract the mean and square the result

=(-29.7%-0.10052)^2+(16.8%-0.10052)^2 +(36.36%-0.10052)^2+(3.9%-0.10052)^2+(22.9%-0.10052)^2

=0.15802215+0.00455355+0.069211086+0.00378471+0.01650711
=0.252078606
Variance=0.252078606/5
0.050415721 or 5.041572% (Rounded to 6 decimal places)

So, the variance is 5.041572%